Eliana wants to make a diorama or an art contest contest. SHE starts with a rectangular sheet of a cardboard box is 12 inches long and 8 inches wide. SHE PLANS to make the diorama box baby cute cutting identical squares from the corners and folding the sides to create a box with an open frontan open box front.
without graphing v(x)=(12-2x)(8-2x)x, describe the end behavior of the graph of the graph of the function y=v(x). Explain .

Answer by stanbon(75887)

Eliana wants to make a diorama or an art contest contest. SHE starts with a rectangular sheet of a cardboard box is 12 inches long and 8 inches wide. SHE PLANS to make the diorama box baby cute cutting identical squares from the corners and folding the sides to create a box with an open frontan open box front.
without graphing v(x)=(12-2x)(8-2x)x, describe the end behavior of the graph of the graph of the function y=v(x). Explain .
-----
Since the highest power term is x*(-2x)(-2x) = 4x^3,
as x goes to + infinity, y will go to + infinity
and as x goes to -infinity, y will go to -infinity.
----------
Cheers,
Stan H.
--------